# Exploring Power Side-Channel Challenges in Embedded Systems Security

### v1

@@ 5.2 Probing
Figure 5: The fingerprint for AES128 in presence of peripheral peaks. The 10 round are detectable in the extracted fingerprint.
A good probing point should be as close as possible to the main processor to avoid having extra noise from other components. However, this is not possible in an off-the-shelf embedded board with lots of components and peripherals. This challenge is addressed by eliminating peripherals and having just a single processor in CW board. This solution seems efficient in applications where firmware binary image is available and has no peripheral interaction. So, the attacker can load it on the CW and apply the PSC attack. However, firmware is usually not available for most of the embedded targets, especially in black-box scenarios. To solve this issue, we propose using a custom-designed interposer with proper considerations for PSC measurements. An interposer is an electrical device acting as an interface between two connections and allows for a third connection to monitor the main signal. For high-speed signals, the interposer handles the impedance matching to avoid signal disturbance. Here, the designed interposer should add extra components for proper power sampling. It should cut the power lines and put a shunt resistor on them. Also, a Low-Pass Filter (LPF) before the shunt resistor is desired to have a stable and denoised VCC (this LPF will remove the noise coming from existing components on the board while the high-frequency information for our algorithm is still there). Then, the main processor should be desoldered to put the interposer on the main PCB and then solder the target processor on the interposer. Although this may require a single interposer for each package type, this can solve the PSC measurement problem for real-world embedded devices.
@@ 5.4 Signal Processing
Figure 6: CFG fingerprinting for a ini parser. All inputs have same execution time while the fingerprint is different.
In this section, we evaluate our proposed frequency-domain fingerprinting technique from a signal-processing perspective. Traditional pattern-matching methods are not robust to time-shift, which are common in embedded firmware running on an OS, and addressing this issue requires additional signal processing to locate the correct position in the trace. In contrast, our method does not require matching the entire power trace, as the relevant information is already contained within the fingerprint. Figure 6 shows the fingerprints for an ini file parser [47]. While all inputs result in the same execution time, the fingerprints differ, indicating different execution paths (potentially new basic blocks). A similarity metric such as Mean Square Error (MSE) or the correlation coefficient can easily detect these differences. The power traces were collected from an STM32F3 microcontroller, running with an 8 MHz clock and no peripheral interaction, at a sampling rate of 50 MS/s, allowing real-time processing.
@@ 5.5 Run-time Overhead
Existing PSC-based methods are unsuitable for real-time use due to their reliance on multiple traces to improve the signal-to-noise ratio. In contrast, wavelet-based fingerprinting can extract accurate fingerprints from a single measurement without noise reduction techniques. Although false positives are inevitable in black-box scenarios, they remain below 30% (A shown in Table III), making our method viable for real-time applications like IDS and fuzzing.
@@ 5.6 Crash Detection and Crash Clustering
Figure 7: Power pattern for hard-faults of a STM32F3 and their comparison.
@@
We trigger the faults in the middle of a function, executing a mock workload to simulate the fault occurring during computation. This will lead the processor to execute a handler for the fault. Fingerprinting the fault handler itself might yield data that could be used to detect or cluster crashes, as different types of failures might cause differences in control flow in the fault handler. However, we observe that the firmware of embedded devices often does not contain error-handling logic in the fault handlers. This is probably because one of these faults occurring is a sign that something went wrong in an unexpected way necessitating a reset of the device. The default fault handler in projects set up using the STM32CubeIDE [48] is just a spinning loop. For the use case of crash detection, detecting this spinning loop is sufficient.
The behavior and, thus, power trace of the hardware after encountering a fault but before executing the software fault handler, however, differs depending on the type of fault that occurs. Figure 7 shows the power pattern for three hard-faults namely Illegal Function Call (function pointer to an invalid address), Illegal Memory Read, and Unaligned Access. The power patterns are distinct for each fault. Depending on the fault, a large peak may be a rising or falling, while the pattern preceding the fault remains the same across all three cases. Note that these power patterns correspond to the hardware handling of the fault before triggering the software handler.
When applied to software testing, this detectable behavioral difference can be used to cluster crashes by the type of fault they cause. For example, if an input triggers a crash due to a function call at an invalid address, and modifying part of that input results in a crash caused by an invalid instruction, we can infer that the crashing input likely corrupts a function pointer. This clustering helps in identifying the root causes of crashes, enabling debugging and vulnerability identification.
@@ 6.1 Real-World Target Fingerprinting
To evaluate the efficiency of the proposed wavelet-based fingerprinting method on a real-world target, we conducted a gray-box test using the UBlox GNSS module without prior knowledge of its firmware or hardware. As shown in Figure 12 in the Appendix 8.1, the module has a shield to reduce noise in the RF section. We removed the shield to access the main processor pins for a closer probing point. The figure illustrates two identical microprocessors without shunt resistors. We then probe the power pins of the second processor, as highlighted in the figure, since the RF signals are routed to it after passing through a mixer. We pass three different commands (namely UBX-CFG-PRT, UBX-CFG-INF and UBX-CFG-DGNSS) to the module and record the power traces.
One of the power traces for these commands is shown in Figure 8 (a). As illustrated, the peripheral peaks are easily detectable, and low-frequency noise is also observable, which complicates fingerprinting. This noise is caused by the other components on the board. Fingerprinting this signal will not produce robust results as the noise highly affects the wavelet transform. However, recalling the wavelet-based fingerprinting subsection, we know that the PC-related information is encoded in high frequencies. Therefore, a high-pass filter removes the low-frequency noise without compromising the desired information. We use the internal high-pass filters of the Tektronix Oscilloscope, and the resulting signal is shown in Figure 8 (b).
Figure 9: The extracted fingerprints for three different commands from a UBlox module
With the low-frequency noise removed, fingerprinting the power traces becomes feasible. As previously mentioned, peripheral peaks can disturb the fingerprinting algorithm. However, they do not contain information relevant to our analysis, as they occur when the processor sends or receives data. Instead, our focus is on the processes applied to the received packets. Therefore, the power trace is clipped, and only the section between the sending and receiving peripheral peaks is used for fingerprinting. The resulting fingerprints are shown in Figure 9. Each command is repeated twice and fingerprinted to show the algorithm’s robustness. The commands are clearly distinguishable, confirming the efficiency of our method.
@@ 6.2 Peripheral Interaction
In this subsection, we evaluate the peripheral interaction detection on a real-world target. High peaks in power consumption are observed during signal transitions on an external interface, where the significant current is either consumed or released. According to Ohm’s law, during a rising edge, a high current is consumed briefly, leading to a voltage drop in the power trace (assuming constant resistance). $\bm{R}=\frac{\bm{V}}{\bm{I}}$ Where R is the resistance, V is the voltage and I is the current. During a falling edge, the current previously consumed is released, increasing the observed voltage. Using this algorithm, we can detect both rising and falling edges on the data lines, enabling the decoding of data from the power trace.
We use the satellite board to evaluate our algorithm in a real-world scenario. The board includes an FRAM, and our goal is to decode the data transferred between the processor and the FRAM. For this experiment, we set the sampling frequency to 1.25 Gs/s. Figure 10 (a) shows the power trace captured by probing one of the board’s power lines with no hardware manipulation (e.g., adding a shunt resistor). As shown in the figure, the high peaks indicate peripheral interactions. Cutting the power trace in the interaction area, we observe a group of peaks (Figure 10 (b)). Zooming on one of the peaks, we see a signal with eight large peaks and smaller peaks in between (Figure 10 (c)). Each peak follows the same pattern. This pattern belongs to the Serial Peripheral Interface (SPI), as eight repeated high peaks represent the clock cycles for a byte transfer. Figure 10 (d) shows the decoded data signal using the abovementioned algorithm. In SPI, data sampling by the receiver is synchronized with the rising edge of the clock (green peaks), meaning that data transitions occur on the falling edge of the clock (red peaks). Classifying the red peaks into three groups allows us to distinguish rising edges, falling edges, and those without transitions. As a result, the data is decoded, and the corresponding binary values are shown.
6.3 Software Composition Analysis
@@
In this section, we evaluate the proposed wavelet-based algorithm to detect a AES algorithm in a real-world target with a real-world embedded RTOS running on it. As a target board, we choose the Nucleo board mentioned in Section 4.3. To simulate a realistic application, we select the Secure MQTT Sensor/Actuator [49] sample from the Zephyr OS project [50] and build it for the Nucleo board without modification.
Without having any customization for PSC measurement or any injected trigger signals, we record the power trace of the processor for a period of two seconds. For the application that we have, it should be more than enough to have a successful communication. The sampling frequency of the oscilloscope is set to 25 MS/s, and the probing point is a power pin on the main processor.
Figure 11 (a) shows the Raw power trace recorded from the target board. The effect of noise is observable on the signal as no data is detectable from the raw power trace. Same as in the previous subsection, we apply a high-pass filter to reduce low-frequency noise. The denoised signal is plotted in Figure 11 (b), and the encryption parts that consume more power are distinguishable in the filtered trace. In the last step, we apply our fingerprinting algorithm on the filtered signal and plot the resulting fingerprint in Figure 11 (c). The rounds of AES encryption are visible and plotted inside red circles in the figure. As the encryption is AES 256 here, it has more rounds than the AES 128 used in Section 5.
Detecting the AES algorithm in a real-world target shows that the proposed fingerprinting algorithm is able to perform across different development boards with different applications.
@@ 7 Discussion
In this paper, we proposed a non-profiling method to eliminate dependencies on specific hardware or firmware. Although signal processing algorithms may introduce performance bottlenecks, this can be mitigated or resolved with more optimized hardware. In our prototype, we used an oscilloscope for sampling and transferred the power traces to our host machine in real-time via a direct Ethernet connection. The wavelet transform was optimized to run in parallel on an Nvidia 3070 Mobile GPU. The time required for data transfer and processing is dependent on the size of the power trace. However, using an oscilloscope connected to a host machine in a real-time scenario can limit processing speed, particularly with large power traces. We aim to address this limitation in our future work by implementing our algorithm on an FPGA that could be a more viable solution. FPGA implementation would eliminate the need to transfer large power traces to an external host, thereby enabling real-time processing and significantly enhancing the speed and efficiency of our algorithm.
@@ 8 Conclusion
In this paper, we systematically examined the application of PSC attacks in the embedded system security context. While PSC attacks have successfully extracted cryptographic keys, expanding their use to broader embedded system applications poses unique challenges. Our analysis identified key factors that reduce the effectiveness of traditional PSC techniques in real-world settings. To address these, we introduced a novel signal processing method that addresses barriers in PSC applications, particularly by eliminating the need for profiling and a shunt resistor. Our experiments on real-world embedded devices demonstrate that this method is both effective and scalable, even with limited knowledge of the target device and firmware, without the need for any hardware modification. Our findings suggest that PSC attacks can be more widely applied in embedded system security by overcoming the discussed challenges. Future work can focus on refining these methods for broader applications and setups, ensuring that PSC becomes a critical tool for embedded systems security.
@@ 8.1 UBlox GNSS Receiver
8.2 Confusion Matrix for Covariate Shift
(a) Train and attack on data from the same distribution (Acc=81%)
(b) Train on random data and attack on AES (Acc=41%)
Figure 13: Effect of covariate shift in instruction disassembly.
Figure 13 shows the confusion matrix mentioned in the Section 4.6.

### v2

@@ 5.2 Probing
Figure 5: CFG fingerprinting for an ini parser. All inputs have the same execution time with different fingerprints.
A good probing point should be as close as possible to the main processor to avoid having extra noise from other components. However, this is impossible in an off-the-shelf device with many components and peripherals. This challenge is addressed by eliminating peripherals and having a single processor in the CW board. This solution seems efficient in applications where the firmware binary image without peripheral interaction is available. So, the attacker can load it on the CW and apply the PSC attack. However, firmware is usually unavailable for most embedded targets, especially in black-box scenarios. To solve this issue, we propose using a custom-designed interposer with considerations for PSC measurements. An interposer is an electrical device acting as an interface between two connections, allowing a third connection to monitor the main signal. For high-speed signals, the interposer handles the impedance matching to avoid signal disturbance. Here, the designed interposer should cut the power lines and put a shunt resistor in place for proper power sampling. Then, the main processor should be desoldered to put the interposer on the main PCB and then solder the target processor on the interposer. Although this may require a single interposer for each package type, this can solve the PSC measurement problem for real-world embedded devices.
@@ 5.3 Noise
Figure 6: Power pattern for hard-faults of a STM32F3 and their comparison (Jitter is removed by a low-pass filter).
@@ 5.4 Signal Processing
This section evaluates our proposed frequency-domain fingerprinting technique from a signal-processing perspective. Traditional pattern-matching methods are not robust to time-shift, which are common in embedded firmware running on an RTOS, and addressing this issue requires additional signal processing to locate the correct position in the trace. In contrast, our method does not require matching the entire power trace, as the relevant information is already contained within the fingerprint. Figure 5 shows the fingerprints for an ini file parser [47]. While all inputs result in the same execution time, the fingerprints differ, indicating different execution paths (potentially new basic blocks). A similarity metric such as Mean Square Error (MSE) or the correlation coefficient can easily detect these differences. The power traces were collected from an STM32F3 microcontroller, running with an 8 MHz clock and no peripheral interaction, at a sampling rate of 50 MS/s, allowing real-time processing.
@@ 5.5 Run-time Overhead
Existing PSC-based methods are unsuitable for real-time use due to their reliance on multiple traces to improve the SNR. In contrast, wavelet-based fingerprinting can extract accurate fingerprints from a single measurement without noise reduction techniques. Although false positives are inevitable in black-box scenarios, they remain below 12% (As shown in Table III), making our method viable for real-time applications like IDS and fuzzing.
@@ 5.6 Crash Detection and Crash Clustering
(a) AVR Fingerprints
(b) ARM Fingerprints
Figure 7: Cross Architecture Fingerprints
@@
We trigger the faults in the middle of a function, executing a mock workload to simulate the fault occurring during computation. This leads the processor to execute a handler for the fault. Fingerprinting the fault handler itself might yield data that could be used to detect or cluster crashes, as different types of failures might cause differences in control flow in the fault handler. However, we observe that the firmware of embedded devices often does not contain error-handling logic in the fault handlers. This is probably because one of these faults occurring is a sign that something went wrong in an unexpected way necessitating a reset of the device. The default fault handler in projects setup using the STM32CubeIDE [48] is just a spinning loop. For the use case of crash detection, detecting this spinning loop is sufficient. The behavior and, thus, power trace of the hardware before executing the software fault handler differs depending on the fault type. Figure 6 shows the power pattern for three hard-faults namely Illegal Function Call (function pointer to an invalid address), Illegal Memory Read, and Unaligned Access. The power patterns are distinct for each fault. Depending on the fault, a large peak may be a rising or falling, while the pattern preceding the fault remains the same across all three cases. Note that these power patterns correspond to the fault’s hardware handling before triggering the software handler.
When applied to software testing, this detectable behavioral difference can be used to cluster crashes by the type of fault they cause. For example, if an input triggers a crash due to a function call at an invalid address, and modifying part of that input results in a crash caused by an invalid instruction, we can infer that the crashing input likely corrupts a function pointer. This clustering helps identify the root causes of crashes, enabling debugging and vulnerability identification.
5.7 Cross Architecture
As explained previously, cross-architecture generalization is a bottleneck for PSC applications. To show that our method can overcome this challenge, we implement the three programs from subsection 5.1 on a different architecture. We choose the ATXMEGA128, an AVR microcontroller with a soldered shunt resistor on the PCB. To run our experiment, we set the sampling frequency of the oscilloscope to 62.5 MS/s and probe the power on the soldered shunt resistor on the PCB. Figure 7 shows the extracted fingerprints and the different execution paths for different programs, while the fingerprints are similar to the ones that we executed on the ARM architecture. This proves the cross-architecture generalizability of our method and the uniqueness of the extracted fingerprint regardless of the underlying hardware.
@@ 6.1 Real-World Target Fingerprinting
To evaluate the efficiency of the proposed wavelet-based fingerprinting method on a real-world target, we conducted a gray-box test using the UBlox GNSS module without prior knowledge of its firmware or hardware. As shown in Figure 16 in the Appendix 8.4, the module has a RF noise-stop shield. We removed the shield to access the main processor pins for a closer probing point. The figure illustrates two identical microprocessors without shunt resistors. We then probe the power pins of the second processor, as highlighted in the figure, since the RF signals are routed to it after passing through a mixer. We pass the UBX-CFG-PRT with different payloads.
Figure 8 (a) shows one of the power traces for these commands. As illustrated, the peripheral peaks are detectable, and low-frequency noise is also observable, complicating fingerprinting. The other components on the board cause this noise. Fingerprinting this signal will not produce robust results as the noise highly affects the wavelet transform. However, recalling the wavelet-based fingerprinting subsection, we know that the PC-related information is encoded in high frequencies. Thus, a high-pass filter removes the low-frequency noise without compromising the desired information. We use a Butterworth high-pass filter of degree five, and the resulting signal is shown in Figure 8 (b).
Figure 9: The extracted fingerprints for different commands from a UBlox module
With the low-frequency noise removed, fingerprinting the power traces becomes feasible. As previously mentioned, peripheral peaks can disturb the fingerprinting algorithm. However, they do not contain information relevant to our analysis, as they occur when the processor sends or receives data. Instead, our focus is on the processes applied to the received packets. Therefore, the power trace is clipped, and only the section between the sending and receiving peripheral peaks is used for fingerprinting. Figure 9 shows the fingerprints, where the differences in the execution path of rejected and accepted payloads are distinguishable.
@@ 6.2 Peripheral Interaction
In this subsection, we evaluate the peripheral interaction detection on a real-world target. High peaks in power consumption are observable during signal transitions on an external interface, where the significant current is either consumed or released. According to Ohm’s law ($\bm{R}=\frac{\bm{V}}{\bm{I}}$, where R is the resistance, V is the voltage, and I is the current), during a rising edge, a high current is consumed briefly, leading to a voltage drop in the power trace (assuming constant resistance). During a falling edge, the current previously consumed is released, increasing the observed voltage. Using this algorithm, we detect both rising and falling edges on the data lines, enabling the decoding of data from the power trace.
We use the satellite board to evaluate our algorithm in a real-world scenario. The board includes a Ferroelectric RAM (FRAM), and our goal is to decode the data transferred between the processor and the FRAM. For this experiment, we set the sampling frequency to 1.25 Gs/s. Figure 10 (a) shows the power trace captured by probing the board’s power lines without hardware manipulation. As shown in the figure, the high peaks indicate peripheral interactions. Cutting the power trace in the interaction area, we observe a group of peaks (Figure 10 (b)). Zooming on one of the peaks, we see a signal with eight large peaks and smaller peaks in between (Figure 10 (c)). This pattern belongs to the Serial Peripheral Interface (SPI), as eight repeated high peaks represent the clock cycles for a byte transfer. Figure 10 (d) shows the decoded data using the abovementioned algorithm. In SPI, data sampling by the receiver is synchronized with the rising edge of the clock (green peaks), meaning that data transitions occur on the falling edge of the clock (red peaks). By classifying the red peaks into three groups, we distinguish rising edges, falling edges, and those without transitions.
6.3 Software Composition Analysis
In this section, we evaluate the proposed wavelet-based algorithm to detect a AES algorithm in real-world targets with real-world embedded RTOS running on them. We designed two experiments namely low-end and high-end.
6.3.1 Low-End
As a low-end target board, we choose the Nucleo board mentioned in Section 4.3. To simulate a realistic application, we select the Secure MQTT Sensor/Actuator [49] sample from the Zephyr OS project [50] and build it for the Nucleo board without modification.
Without customization for PSC measurement or any injected trigger signals, we record the power trace of the processor for two seconds. For our application, it should be more than enough to have a successful communication. The sampling frequency of the oscilloscope is set to 25 MS/s, and the probing point is a power pin on the main processor.
Figure 11 (a) shows the Raw power trace recorded from the target board. The effect of noise is observable on the signal as no data is detectable from the raw power trace. Same as in the previous subsection, we apply a high-pass filter to reduce low-frequency noise. Figure 11 (b) shows the denoised signal and the encryption parts that consume more power are distinguishable in the filtered trace. In the last step, we apply our fingerprinting algorithm on the filtered signal and plot the resulting fingerprint in Figure 11 (c). The rounds of AES encryption are visible and plotted inside red circles in the figure. As the encryption is AES 256 here, it has more rounds than the AES 128 used in Section 5.
6.3.2 High-End
Figure 12: Fingerprint of an AES running on the Raspberry Pi Zero 2 W.
As a high-end target, we choose the Raspberry Pi Zero running the Raspbian kernel on it. The sampling frequency of the oscilloscope is set to 250 MS/s. We establish a TLS connection on it and measure the power trace during execution. After applying a low-pass filter, we apply the fingerprinting algorithm. Figure 12 shows the extracted fingerprint, and the AES rounds are again visible in this plot.
Detecting the AES algorithm in a real-world target shows that the proposed fingerprinting algorithm is able to perform across different development boards with different architectures.
@@ 7 Discussion
In this paper, we proposed a fingerprinting method to eliminate dependencies on specific hardware or firmware. Although signal processing algorithms may introduce performance bottlenecks, this can be mitigated or resolved with more optimized hardware. In our prototype, we used an oscilloscope for sampling and transferred the power traces to our host machine in real-time via a direct Ethernet connection. The wavelet transform was optimized to run in parallel on an Nvidia 3070 Mobile GPU. The time required for data transfer and processing is dependent on the size of the power trace. However, using an oscilloscope connected to a host machine in a real-time scenario can limit processing speed, particularly with large power traces. We aim to address this limitation in our future work by implementing our algorithm on an FPGA that could be a more viable solution. FPGA implementation would eliminate the need to transfer large power traces to an external host, thereby enabling real-time processing and significantly enhancing the speed and efficiency of our algorithm.
@@ 8 Conclusion
In this paper, we systematically examined the application of PSC attacks in the embedded system security context. While PSC attacks have successfully extracted cryptographic keys, expanding their use to broader embedded system applications poses unique challenges. Our analysis identified key factors that reduce the effectiveness of traditional PSC techniques in real-world settings. To address these barriers in PSC applications, we introduced a novel signal processing method, particularly by eliminating the need for profiling the same target and a shunt resistor. Our experiments on real-world embedded devices demonstrate that this method is both effective and scalable, even with limited knowledge of the target device and firmware, without the need for any hardware modification. Our findings suggest that PSC attacks can be more widely applied in embedded system security by overcoming the discussed challenges. Future work can focus on refining these methods for broader applications and setups, ensuring that PSC becomes a critical tool for embedded systems security.
@@
8.1 Experimental Setup
Figure 13 shows our experimental setup for the Raspberry Pi experiment. As mentioned before, we use a Tektronix MSO66B oscilloscope for power measurement.
Figure 13: Our Experimental setup
8.2 Confusion matrix for different probing points
Figure 14 shows the confusion matrix for two probing points. As illustrated in the figure, the accuracy drops drastically when the probing point is on the main power supply of the board.
(a) Probing on processor’s pins (ACC=91%)
(b) Probing on board’s pins (ACC=53%)
8.3 Confusion Matrix for Covariate Shift
Figure 15 shows the confusion matrix mentioned in the Section 4.6.
(a) Train and attack on data from the same distribution (Acc=81%)
(b) Train on random data and attack on AES (Acc=41%)
Figure 15: Effect of covariate shift in instruction disassembly.